Circular No. 8091

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003bt Further to IAUC 8090, M. Schwartz and P. R. Holvorcem report the

LOTOSS (cf. IAUC 7906) discovery of a supernova (mag 16.7) on unfiltered Tenagra II CCD images taken on Mar. 9.3 and 10.3 UT. SN 2003bt is located at α = 10^{h}55^{m}22^{s}.48, δ = −7^{o}39′12′′.3 (equinox 2000.0), which is 11′′.9 west and 0′′.1 south of the host galaxy’s nucleus. An image taken on Feb. 2.4 showed nothing at this position (limiting mag ∼ 18.5). T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003bt, obtained by M. Calkins on Mar. 11.25 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova, ∼ 2 weeks past maximum.

SUPERNOVAE 2003bm, 2003bp, AND 2003br Matheson et al. also write that a spectrum of SN 2003bm (cf. IAUC

8086), obtained as above by P. Berlind on Mar. 7.21 UT, shows it to be a type-Ic supernova. A spectrum of SN 2003bp (cf. IAUC 8089), obtained by Calkins on Mar. 10.23, shows it to be a type-Ib supernova. The helium features are relatively weak. This may be a transitional object between types Ib and Ic. A spectrum (range 370-750 nm) of SN 2003br (cf. IAUC 8090), obtained by Calkins on Mar. 11.47, shows it to be a type-II supernova.

P. Nugent and D. A. Howell, Lawrence Berkeley National Laboratory, report that a spectrum (range 360–900 nm) of SN 2003bp, obtained on Mar. 8.22 UT with the Kitt Peak 2.1-m reflector (+ GCAM), shows it to be an early type-Ib supernova, consistent with the photometry (cf. IAUC 8089). SN 2003bp clearly shows He I lines at 587.6, 667.8 and 706.5 nm. The He I lines are somewhat weaker than those of SNe 1998dt and 1999dn (see Matheson et al. 2001, A.J. 121, 1648).

CATALOGUE OF COMETARY ORBITS 2003 A new edition of the Catalogue of Cometary Orbits, with entries com-

plete through February, is now available. The 169-page Catalogue sells for $40.00 ($60.00 for airmail delivery outside North America). There is also an e-mail edition (with just the basic orbital information) for $75.00. Orders should be sent to Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A., by FAX to +1-617-495-7231 or by e-mail to mlohmiller@cfa.harvard.edu. Payment can be by credit card or by check made payable to Smithsonian Astrophysical Observatory.

2003 March 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green